Pearl Products Limited of Shenzhen, China, manufactures and distributes toys throughout South East Asia. Three cubic centimeters (cc) of solvent H300 are required to manufacture each unit of Supermix, one of the company’s products. The company is now planning raw materials needs for the third quarter, the quarter in which peak sales of Supermix occur. To keep production and sales moving smoothly, the company has the following inventory requirements:

a. The finished goods inventory on hand at the end of each month must be equal to 4,000 units of Supermix plus 25% of the next month’s sales. The finished goods inventory on June 30 is budgeted to be 21,250 units.

b. The raw materials inventory on hand at the end of each month must be equal to one-half of the following month’s production needs for raw materials. The raw materials inventory on June 30 is budgeted to be 88,000 cc of solvent H300.

c. The company maintains no work in process inventories.

A sales budget for Supermix for the last six months of the year follows.

Budgeted Sales
in Units
July 69,000
August 74,000
September 84,000
October 64,000
November 54,000
December 44,000

Required:

1. Prepare a production budget for Supermix for the months July, August, September, and October.


3. Prepare a direct materials budget showing the quantity of solvent H300 to be purchased for July, August, and September, and for the quarter in total.

Please show your work
Answer #1
Required 1 :
Pearl Products Limited
Production Budget
July August September October
Budgeted unit sales 69000 74000 84000 64000
(+) Desired units of ending finished goods inventory ( 4000 + 25% of next month's sales ) 22500 25000 20000 17500
Total needs 91500 99000 104000 81500
(-) Units of Beginning finished goods inventory 21250 22500 25000 20000
Required production in units 70250 76500 79000 61500
Note : October's ending inventory = 4000 + ( 25% * November unit sales ) = 4000 + ( 25% * 54000 ) 17500
Required 2 :
Preal Products Limited
Direct materials budget
July August September Third Quarter
Required production in units 70250 76500 79000 225750
Units of raw materials needed per unit of finished goods 3 cc 3 cc 3 cc 3 cc
Units of raw materials needed to meet production 210750 229500 237000 677250
(+) Desired units of ending raw materials inventory ( One-half of following month's production needs ) 114750 118500 92250 92250
Total units of raw materials needed 325500 348000 329250 769500
(-) Units of Beginning raw materials inventory 88000 114750 118500 88000
Units of raw materials to be purchased 237500 233250 210750 681500
Note : September's ending inventory = 1/2 * October month's production needs = 1/2 * ( 61500 * 3 ) 92250